OpenStudy (anonymous):

if the sum of a,b and c is twice the sum of a minus b and a minus c, then a= ?

OpenStudy (jamesj):

So a+b+c = 2(a-b-c) = 2a - 2b - 2c Now solve for a

OpenStudy (anonymous):

a+b+c=2a-2b-2c -a=-3b-3c a=3b+3c